# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ 5 Constant-Roll Kinetically Driven Inflation
At this point, we constrain the model using constant-roll conditions (3.7), (3.9), and (3.13). Using relation $Q=P+V=X+\frac{\gamma}{2}X^{2}$ we can write
@@
HCR: $\frac{\ddot{H}}{2H\dot{H}}=\beta_{\text{\tiny{H}}}\Rightarrow\frac{\ddot{\phi}\left(1+\gamma\dot{\phi}^{2}\right)}{H\dot{\phi}\left(1+\frac{1}{2}\gamma\dot{\phi}^{2}\right)}=\beta_{\text{\tiny{H}}},$ (5.7)
PCR: $\frac{\ddot{\phi}}{H\dot{\phi}}=c_{s}^{2}\beta_{\text{\tiny{V}}}-\frac{c_{s}^{2}\left(\dot{\phi}^{2}Q_{,X\phi}-Q_{\phi}\right)}{H\dot{\phi}\,Q_{,X}}\Rightarrow\frac{\ddot{\phi}\left(1+\frac{3}{2}\gamma\dot{\phi}^{2}\right)}{H\dot{\phi}\left(1+\frac{1}{2}\gamma\dot{\phi}^{2}\right)}=\beta_{\text{\tiny{V}}},$ (5.8)
where we have used equation (5.3) to eliminate $\dot{H}$ and $\ddot{H}$ from equations. Since three constant-roll conditions are different, we expect to find three different sets of solutions. However, when these conditions approach $\gamma\dot{\phi}^{2}\gg 1$ and $\gamma\dot{\phi}^{2}\ll 1$ limits, they become similar, which we will discuss in detail in the next subsection.
5.1 Quasi-Canonical Phase
In the limit of $\gamma\dot{\phi}^{2}\ll 1$ the constant $k$ inflation (5.1) reduces to the canonical case $P(X,\phi)\approx X-V(\phi)$, while in the limit of $\gamma\dot{\phi}^{2}\gg 1$ it takes the following form:
$P(X,\phi)\approx\frac{\gamma}{2}X^{2}-V(\phi).$ (5.9)
In the canonical limit $\gamma\dot{\phi}^{2}\ll 1$, three constant-roll conditions (5.6), (5.7), and (5.8) become identical and we have $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}$. On the other hand, in $\gamma\dot{\phi}^{2}\gg 1$ limit, when relation (5.9) holds, our constant-roll conditions become
$\frac{\ddot{\phi}}{H\dot{\phi}}=\beta_{\text{\tiny{F}}},\quad\quad 2\frac{\ddot{\phi}}{H\dot{\phi}}=\beta_{\text{\tiny{H}}},\quad\quad 3\frac{\ddot{\phi}}{H\dot{\phi}}=\beta_{\text{\tiny{V}}}.$ (5.10)
The three constant-roll conditions can be seen to be similar again when $6\beta_{\text{\tiny{F}}}=3\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{V}}}$. This behavior was already expected. Comparing (5.9) with (4.1), we realize that in $\gamma\dot{\phi}^{2}\gg 1$ limit, $k$ inflation reduces to a constant propagation model with $c_{s}=1/\sqrt{3}$ and $\mathcal{F}=\gamma/2$, and we anticipate encountering the same quasi-canonical behavior observed in that context. As we discussed in section 4, in a constant propagation model with constant $\mathcal{F}$ three constant-roll conditions are analogous (see equation (4.7) and the discussion in the next paragraph). It is also noteworthy that according to the definition of the PCR condition, USR occurs at $\beta_{\text{\tiny{V}}}=-3$, which in (5.9) case corresponds to $\beta_{\text{\tiny{H}}}=-2$ and $\beta_{\text{\tiny{F}}}=-1$ (compare it with the canonical case that USR happens at $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}=-3$).
We recall for in the case of FCR condition $\dot{\phi}$ grows rapidly for $\beta_{\text{\tiny{F}}}>0$, while it decays rapidly for $\beta_{\text{\tiny{F}}}<0$. In $k$ inflation, $\gamma\dot{\phi}^{2}$ can become much larger than 1 at the early stages of inflation by choosing a large $\gamma$. At this stage, the approximation (5.9) will be viable. If $\beta_{\text{\tiny{F}}}$ is negative, as $\dot{\phi}$ decays and $\gamma$ is constant, the first term in (5.1) will eventually surpass the second term, leading to the canonical case $P(X,\phi)\approx X-V(\phi)$. In other words, our model naturally transits into the canonical case. Obviously, if $\gamma$ is too large or $\beta_{\text{\tiny{F}}}$ is too small (which results in a slow decaying speed), there may not be enough time for the model to transit into the canonical case during inflation. The situation is opposite when $\beta_{\text{\tiny{F}}}$ is positive and $\dot{\phi}$ is growing. In this case, since $X$ is growing rapidly, the second term in equation (5.1) will finally become dominant. That means we can start the inflation with canonical case and then move on to (5.9) case.
In the limit of $\gamma\dot{\phi}^{2}\gg 1$, equations (5.2) and (5.3) are simplified as
$3H^{2}=\frac{3}{8}\gamma H^{4}\phi_{,N}^{4}+V,$ (5.11)
$H_{,N}=-\frac{1}{4}\gamma H^{3}\phi_{,N}^{4},$ (5.12)
where we have written all derivatives in terms of $N$. Since this model is a special case of constant propagation model (4.1), all results derived in section 4 apply here as well. That means we can use Hubble parameters (2.13), (2.14), (2.15), and (2.29), potentials (4.18),(4.19),(4.20), and (4.21), slow-roll parameters (2.19), (2.23), (2.27), and (2.34), and spectral indexes (2.35), (2.36), and (2.37). What we were unable to derive in section 4, was the field and accordingly $V(\phi)$. For solutions (2.13), (2.14), (2.15), and (2.29) we can integrate equations (5.11) or (5.12) to find the following solutions for the field respectively:
$\phi(N)=\pm\left(-\frac{64}{\gamma M^{2}\beta_{\text{\tiny{H}}}^{3}}\right)^{1/4}e^{\frac{1}{2}\beta_{\text{\tiny{H}}}N}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-e^{2\beta_{\text{\tiny{H}}}N}\right)+\phi_{c},$ (5.13)
$\phi(N)=\pm\left(\frac{64}{\gamma M^{2}\beta_{\text{\tiny{H}}}^{3}}\right)^{1/4}e^{\frac{1}{2}\beta_{\text{\tiny{H}}}N}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};e^{2\beta_{\text{\tiny{H}}}N}\right)+\phi_{c},$ (5.14)
$\phi(N)=\pm\left(-\frac{64}{\gamma M^{2}\beta_{\text{\tiny{H}}}^{3}}\right)^{1/4}e^{\frac{1}{2}\beta_{\text{\tiny{H}}}N}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};e^{2\beta_{\text{\tiny{H}}}N}\right)+\phi_{c},$ (5.15)
$\phi(N)=\pm\left(\frac{32}{\gamma M^{2}}\right)^{1/4}\sqrt{-N}+\phi_{c},$ (5.16)
where ${}_{2}F_{1}$ is the hypergeometric function. To find $V(\phi)$ we need to solve the above equations to find $N$ as a function of $\phi$. However, due to the presence of the hypergeometric function in solutions (5.13), (5.14), and (5.15), $N(\phi)$ cannot be derived. For the fourth solution (5.16), $V(\phi)$ can be found as
$V(\phi)=\frac{3}{4}M^{2}\left(\sqrt{\frac{\gamma M^{2}}{2}}(\phi-\phi_{c})^{2}-1\right).$ (5.17)
With all the solutions in hand, we can now discuss which ranges of parameters $\beta_{\text{\tiny{H}}}$ and $N$ lead to inflation. Here the Hubble and slow-roll parameters are identical to those in the canonical case, so allowed ranges of $\beta_{\text{\tiny{H}}}$ and $N$ are also identical unless the field becomes imaginary. In the inflationary range, solutions (5.13), (5.14), and (5.16) are real, but solution (5.15) is imaginary. However, the imaginary part of this solution is constant, so it can always be absorbed by redefining the integration constant as $\phi_{c}\rightarrow\phi_{c}-\Im(\phi(N))$. Therefore, these fields do not further constrain the parameters’ ranges.
Now that we have fields, we can compare $\gamma\dot{\phi}^{2}$ with 1 to see how long the approximation (5.9) remains valid. For solutions (5.13) and (5.15) we can write
$\gamma\dot{\phi}^{2}=\gamma H^{2}\phi_{,N}^{2}=2M\sqrt{\gamma\,|\beta_{\text{\tiny{H}}}|}\,e^{\beta_{\text{\tiny{H}}}\left(N-N_{\text{ini}}\right)},$ (5.18)
where we have assumed that inflation starts at $N=N_{\text{ini}}$. Note that this relation is consistent with the general relation (3.8) since we have $\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{F}}}$ in $\gamma\dot{\phi}^{2}\gg 1$ limit. We now define parameter $\delta$ such that the approximation (5.9) starts to break at the critical point $\gamma\dot{\phi}_{\text{cri}}^{2}=\delta$. Using equation (5.18), we can find the number e-fold at the critical point as
@@ 5.1 Quasi-Canonical Phase
As an example, if $\delta=10,M=1,\gamma=10^{5}$, and $\beta_{\text{\tiny{H}}}=-0.5$ then $N_{\text{cri}}\approx 7.6+N_{\text{ini}}$. Therefore, approximation (5.9) is valid for the first 7.6 e-folds of the inflation. After this point the first and second terms in relation (5.1) are comparable and we must use exact solutions. When $\dot{\phi}$ continues to decay, the model will eventually reach the canonical case.
Note that relation (5.18) can also be used for solution (5.14). However, in this solution $\beta_{\text{\tiny{H}}}$ is positive and $\dot{\phi}$ is growing rather than decaying. Which means inflation can be started from the canonical case and then transit to (5.9) case. To evaluate the number e-fold at the critical point (the moment when canonical approximation starts to break) we use solution (2.20) to write
$\gamma\dot{\phi}^{2}=\gamma H^{2}\phi_{,N}^{2}=2\gamma M^{2}\beta_{\text{\tiny{H}}}e^{2\beta_{\text{\tiny{H}}}(N-N_{\text{ini}})},$ (5.20)
where we have assumed that inflation starts at $N=N_{\text{ini}}$. Again, this condition is consistent with relation (3.8) since in the canonical limit we have $\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{F}}}$. Similar to (5.19), by defining a proper $\delta$ parameter, we can find the number e-fold at the critical point as
@@
For instance, for $\delta=0.1,M=1,\gamma=10^{-5}$, and $\beta_{\text{\tiny{H}}}=0.5$ we find $N_{\text{cri}}\approx 9.2+N_{\text{ini}}$. That means canonical approximation is valid at the beginning of inflation until it breaks at around $9.2$ e-folds. After this point we must use exact solutions. As $\dot{\phi}$ grows, we gradually enter the next phase where we can use approximate relation (5.9). For the last solution (5.16), since $\ddot{\phi}$ is zero and $\dot{\phi}$ constant, there is no transition between (5.9) and canonical case.
5.2 Exact Solutions
We will now solve equations of motion analytically without considering any approximations. Writing time derivatives in equations (5.2) and (5.3) in terms of $N$ we will have
@@ 5.2 Exact Solutions
Let us first derive the Hubble parameter for all three constant-roll conditions (5.6), (5.7), and (5.8). As we discussed in section 3, for the HCR condition it is possible to derive Hubble parameter even without knowing the exact form of the action. As a result, we can again use solutions (2.13), (2.14), (2.15), and (2.29) for Hubble parameters. For the PCR condition (5.8) Hubble parameter cannot be found analytically. For the FCR condition, after writing the time derivatives in (5.6) in terms of $N$ and removing $\phi_{,N}$ and $\phi_{,NN}$ using equation (5.23), we get
@@
Integrating out, we can find the following solution for this equation:
$H^{2}=\frac{1}{2\gamma\beta_{\text{\tiny{F}}}}\left[\tilde{M}^{2}-\left(e^{2\beta_{\text{\tiny{F}}}(N-\tilde{N})}+1\right)^{2}\right],$ (5.25)
where $\tilde{M}$ and $\tilde{N}$ are two complex constants, which similar to our discussion in section 2, they should be chosen properly so that (5.25) leads to a viable inflationary scenario. When $\beta_{\text{\tiny{F}}}$ is exactly zero, we will find the solution (2.29) again. Thus, we identified eight constant-roll solutions. Four solutions from (5.6), three solutions from (5.7), and one solution for $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=0$ case. In order to compare the FCR and HCR conditions, we can proceed to derive other parameters of the model. But, the PCR condition can only be treated numerically. In the next subsection, we investigate a practical inflationary scenario and derive the power spectrum and other relevant parameters numerically to compare all three constant-roll conditions.
5.3 PBH Formation
It is clarified that a sharp enhancement in the amplitude of the power spectrum between small and large scales can be responsible for the production of PBHs [57]. To achieve such an enhancement one needs to look beyond the slow-roll limit and/or canonical case. In constant $k$ model, by choosing a large value for the constant-roll parameter, the slow-roll approximation can easily be violated. Moreover, the model is non-canonical and has variable sound speed, which affect the power enhancement.
To understand the enhancement mechanism of power spectrum we look at the differential equation governing the curvature perturbation,
$u_{k}^{\prime\prime}+2\frac{z^{\prime}}{z}u_{k}^{\prime}+k^{2}u_{k}=0,$ (5.26)
where for $z^{\prime}/z$ we have
$\frac{z^{\prime}}{z}=aH\left(1+\frac{\epsilon_{2}}{2}+q_{1}\right),$ (5.27)
and $q_{1}$ is defined in equation (A.4). Equation (5.26) is similar to the equation of motion of a harmonic oscillator with friction. When $z^{\prime}/z$ becomes negative, the friction term acts as a driving term, resulting in a sharp increase in the power spectrum. In the canonical constant-roll model, when we have $\epsilon_{2}\approx 2\beta$, to achieve negative $z^{\prime}/z$ we need to choose $\beta<-1$. In the non-canonical case with variable sound speed $q_{1}$ is non-zero and should be taken into account.
Now we construct a four-stage constant-roll model, capable of enhancing power spectrum, satisfying CMB constraints, and providing a graceful exit from inflation. First, we discuss the scenario for the canonical case, then we apply it to $k$ inflation under the three constant-roll conditions (5.6), (5.7), and (5.8).
ptRefer to caption    Refer to captionptRefer to caption      Refer to captionptRefer to caption      Refer to caption

### v3

@@ 5.1 Background Dynamics
We constrain the model using constant-roll conditions (3.7), (3.9), and (3.13). Using the relation $Q=P+V=X+\frac{\gamma}{2}X^{2}$, we have
@@
[…]
PCR: $\frac{1}{c_{s}^{2}}\frac{\ddot{\phi}}{H\dot{\phi}}=\beta_{\text{\tiny{V}}},$ (5.11)
where we have used equation (5.3) to eliminate $\dot{H}$ and $\ddot{H}$ from equations. Note that here $\lambda$ and $c_{s}$ are evolving variables and these constant-roll conditions are not homologous in general. From this system of equations, one can see that one important parameter for the evolution is the product $\gamma X$ (or $c_{s}^{2}$). For $\gamma X\gg 1$, we have $P\approx\frac{\gamma}{2}X^{2}-V$, and the kinetic energy is dominated by the non-canonical term, but for $\gamma X\ll 1$, action reduces to the canonical case. In the latter case, the three constant-roll conditions are homologous with $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}$, and the background parameters $\epsilon_{n}$, $c_{s}$, and $q_{n}$ reduce to the canonical values discussed in section 2. For $\gamma X\gg 1$, the sound speed is approximated by $c_{s}^{2}\approx 1/3$ with $q_{1}=0$, which means that no matter which of the constant-roll conditions are used, inflation is in a constant propagation regime with $\mathcal{F}=\gamma/2$ (see (4.1)), and therefore we can get use of the results discussed in section 4. In this regime, since $\mathcal{F}$ is constant, the three models are homologous and the constant-roll parameters are related by $6\beta_{\text{\tiny{F}}}=3\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{V}}}$.
The interesting feature of the model is that it can naturally transition between the two limiting phases. In the case when constant-roll parameter is negative, one can think of an inflationary scenario that starts in $\gamma X\gg 1$ regime and gradually transitions to a canonical constant-roll system. On the contrary, when constant-roll parameter is positive and $X$ is growing in time, the model evolves in the opposite direction, transitioning from the canonical to $\gamma X\gg 1$ phase. Section 5.2 will present an example where we manipulate the constant-roll parameter to illustrate these transition intervals.
While the three constant-roll conditions are homologous during the two limiting phases, the transition between these phases is strongly influenced by the chosen condition. Here we briefly discuss when the constant propagation approximation breaks and we can expect different consequences from different constant-roll constraints. If the constant-roll parameter is negative (cases (2.13) and (2.15)) and we start the inflation in $c_{s}^{2}=1/3$ phase, we can write
$\gamma\dot{\phi}^{2}=2M\sqrt{\gamma\,|\beta_{\text{\tiny{H}}}|}\,e^{\beta_{\text{\tiny{H}}}\left(N-N_{\text{ini}}\right)},$ (5.12)
where we have assumed that inflation starts at $N=N_{\text{ini}}$. Note that this relation is consistent with the general relation (3.8) since we have $\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{F}}}$ in $\gamma\dot{\phi}^{2}\gg 1$ limit. We now define parameter $\delta$ such that the approximation $P\approx\frac{\gamma}{2}X^{2}-V$ starts to break at the critical value $\gamma\dot{\phi}_{\text{cri}}^{2}=\delta$ [^5] It is important to note that the transition between these two phases occurs smoothly. Defining a critical time is merely to demarcate the boundaries of each phase and does not mean a sudden shift from one phase to the other within the model. The critical time serves as a convenient marker to outline the range where our constant propagation approximation becomes less accurate and to highlight the regimes at which constant-roll conditions result in novel consequences and are of particular interest. Using equation (5.12), one can find the critical $e$-folds number as
@@
As an example, if $\delta=10,M=1,\gamma=10^{5}$, and $\beta_{\text{\tiny{H}}}=-0.5$ then $N_{\text{cri}}\approx 7.6+N_{\text{ini}}$. Therefore, approximation $c_{s}^{2}\approx 1/3$ is valid for the first 7.6 $e$-folds of the inflation. After this point the two terms in relation (5.1) are comparable and we must use exact solutions. When $\dot{\phi}$ continues to decay, the model will eventually reach the canonical limit.
Now consider the case when constant-roll parameter is positive (case (2.14)) and model starts in the canonical phase. To evaluate the moment when canonical approximation starts to break, we use solution (2.20) to write
$\gamma\dot{\phi}^{2}=2\gamma M^{2}\beta_{\text{\tiny{H}}}e^{2\beta_{\text{\tiny{H}}}(N-N_{\text{ini}})},$ (5.14)
where we have assumed that inflation starts at $N=N_{\text{ini}}$. Again, this condition is consistent with relation (3.8) since in the canonical limit we have $\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{F}}}$. Similar to (5.13), by defining a proper $\delta$ parameter, we can find the number of $e$-folds that canonical constant-roll lasts by
@@
For instance, for $\delta=0.1,M=1,\gamma=10^{-5}$, and $\beta_{\text{\tiny{H}}}=0.5$ we find $N_{\text{cri}}\approx 9.2+N_{\text{ini}}$. That means canonical approximation is valid at the beginning of inflation until it breaks at around $9.2$ $e$-folds. After this point, we must use exact solutions. As $\dot{\phi}$ grows, we gradually enter the second constant propagation phase with $c_{s}^{2}=1/3$.
Let us focus more on the background dynamics of this model. The time derivatives in equations (5.2) and (5.3) can be written in terms of $N$ to give
@@
In HCR models, the Hubble parameter is the same as the constant-roll canonical case. For FCR models, however, after writing the time derivatives in constant-roll condition (5.9) in terms of $N$ and removing $\phi_{,N}$ and $\phi_{,NN}$, using equation (5.17), we get
@@
Fortunately, this can be integrated to give
$H^{2}=\frac{1}{2\gamma\beta_{\text{\tiny{F}}}}\left[C_{1}-\left(C_{2}e^{2\beta_{\text{\tiny{F}}}N}+1\right)^{2}\right],$ (5.19)
where $C_{1}$ and $C_{2}$ are two complex constants, which similar to our discussion in section 2, they should be chosen properly so that (5.19) leads to a viable inflationary scenario.
In order to compare the FCR and HCR conditions, one can proceed to derive other parameters of the model analytically, though the PCR condition can only be treated numerically. To explore these models and show the details of transitions, we would rather perform a full numerical study on an inflationary scenario consisting of multiple constant-roll stages instead. In the following subsection, we construct a four-stage constant-roll model and study the evolution of scalar perturbations in this scenario for the three constant-roll constraints. The time dependent background parameters in each case can affect the horizon crossing time of modes, and besides other mechanisms, translate into a scale dependent amplification of the power spectrum.
5.2 Evolution of Scalar Perturbations
Having examined the general dynamical properties of the background parameters, we now proceed to investigate scalar perturbations of the model. The curvature perturbations of wavenumber $k$ for the general single field action (A.1) satisfy the following differential equation:
$\mathcal{R}_{k}^{\prime\prime}+2\frac{z^{\prime}}{z}\mathcal{R}_{k}^{\prime}+c_{s}^{2}k^{2}\mathcal{R}_{k}=0,$ (5.20)
where $z\equiv a\sqrt{2\epsilon_{1}}/c_{s}$ and $z^{\prime}/z$ can be expressed in terms of the background parameters as
$\frac{z^{\prime}}{z}=aH\left(1+\frac{\epsilon_{2}}{2}-q_{1}\right).$ (5.21)
A full analysis of equation (5.20) shows the phenomenological consequences of a departure from slow-roll and/or canonical models. One can (plot the bracket in equation (5.21) to) find the ranges of constant-roll parameter at which $z^{\prime}/z$ is negative implying that the friction term (proportional to $z^{\prime}/z$) acts as driving term and the would be decaying solution of (5.21) grows in time. This will contaminate the nearly constant solution $\mathcal{R}_{k}$, leading to an amplified value at late time, for superhorizon scales. It is clarified that such an (amplitude) enhancement can be responsible for the production of PBHs [67].
As alluded to above, in canonical constant-roll models $\epsilon_{2}\approx 2\beta$, hence $\beta<-1$ is the sufficient condition for the mode amplification. In constant-roll models of our interest, $\epsilon_{2}$ and $q_{1}$ are proportional to the $\ddot{\phi}/(H\dot{\phi})$ which is constrained by equations (5.9)-(5.11). To discuss the three models altogether, we define a function as
$\Lambda(c_{s})\equiv\begin{cases}1,&\text{in FCR}\\ \lambda^{-1},&\text{in HCR}\\ c_{s}^{2},&\text{in PCR}\end{cases}.$ (5.22)
With this definition, the constant-roll conditions can be represented in a unified form, $\ddot{\phi}/(H\dot{\phi})=\beta\Lambda$, where $\beta$ can be any of the three constant-roll parameters. Using equations (5.7) and (5.8) we find the general condition to achieve negative $z^{\prime}/z$ as
$\beta<\beta_{\text{cri}}\equiv\frac{2}{(3c_{s}^{2}-5)}\frac{1}{\Lambda}.$ (5.23)
In Figure 10, we have plotted $\beta_{\text{cri}}$ in terms of $c_{s}^{2}$ for each constant-roll condition. In the canonical limit, this condition necessitates $\beta<-1$, while in $\gamma X\gg 1$ limit, it reduces to $\beta<-1/2$, $-1$, and $-3/2$ for FCR, HCR, and PCR conditions, respectively.
$\beta_{\text{cri}}$
$c_{s}^{2}$
Figure 10: The plot of the $\beta_{\text{cri}}$ function defined in equation (5.23) as a function of $c_{s}^{2}$ for the three constant-roll conditions. Selecting a constant-roll parameter smaller than $\beta_{\text{cri}}$ is the sufficient condition for achieving a negative $z^{\prime}/z$, which can lead to mode amplification and a subsequent burst in the power spectrum. The vertical grid line corresponds to the $c_{s}^{2}=1/3$.
Refer to caption     Refer to captionRefer to caption       Refer to caption